OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −4x + y = −2? y equals four x plus c y equals negative four x plus c y equals negative one over four times x plus c y equals one over four times x plus c

OpenStudy (ikram002p):

−4x + y = −2 change it to slope intercet formula y=4x-2 slope = 4

OpenStudy (ikram002p):

so the slope of perpendiculer = -1/slope=-1/4 so perpendicular line :- y=-1/4 x +c

OpenStudy (anonymous):

So, C?

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

yess :D